# BanglaTalk: Towards Real-Time Speech Assistance for Bengali Regional Dialects

### v1

@@ 4.1 Bengali Speech Assistants
Several speech assistant systems have been developed for the Bengali language in recent years. The ALAPI system (Hasan et al., 2021) introduces an open-domain Bengali conversational agent that processes recorded audio queries from users and generates response audio using AI techniques alongside a custom-built database. Shohojogi (Arnab et al., 2023), designed for the banking sector, provides voice-based customer support in Bengali. It integrates Wav2Vec2-based ASR (Baevski et al., 2020), query summarization, Google Text-to-Speech (gTTS), and a doc2vec model to retrieve relevant information for responses. Adrisya Sahayak (Sultan et al., 2021) presents a desktop-based virtual speech assistant for visually impaired Bengali speakers, supporting computer operations, peripheral devices, and home appliance control.
In the healthcare domain, a voice-enabled Artificial Conversational Entity has been developed to automate service interactions in Bengali (Pranto et al., 2021). This system leverages a domain-specific database and similarity-matching strategies to generate responses to user queries. Extending beyond monolingual assistants, Disha (Ullah et al., 2024) is a humanoid virtual assistant capable of interacting in both Bengali and English. It can address financial queries and perform real-time transactions. Beyond general-purpose applications, Bengali voice assistants have also been deployed in specialized domains such as providing information on metro rail services in Bangladesh (Rahman et al., 2024), assisting farmers with agricultural queries (Divakar et al., 2021), and offering accessible feminine healthcare support for marginalized women (Puja et al., 2024).
@@ 4.2 English Speech Assistants
Voice assistants such as Amazon Alexa, Apple Siri, Microsoft Cortana, and Google Assistant have become integral to modern human-computer interaction (López et al., 2017). These systems, powered by artificial intelligence, are increasingly influencing daily life and societal practices (Subhash et al., 2020). Their development has been driven by advances in both signal processing and machine learning, which form the technological foundation of voice-based interaction (Haeb-Umbach et al., 2019).
Beyond their technical design, researchers have examined the broader impact of these systems. For example, Flavián et al. (2023) demonstrated that, compared to text-based recommendations, voice-based recommendations delivered by assistants have a stronger influence on consumer decision-making. User studies also indicate that voice assistants are most frequently employed for activities such as music, information search, and smart home (IoT) control (Ammari et al., 2019). Furthermore, their role has extended into education and healthcare contexts, where they support learning environments and assist older adults in managing everyday activities (Terzopoulos and Satratzemi, 2020; Oewel et al., 2023).
@@ 4.3 Automatic Speech Recognition in Bengali
Research on Automatic Speech Recognition (ASR) in Bengali has been addressed through both system development and survey-driven studies (Mridha et al., 2022; Tasnia et al., 2023; Sultana et al., 2021). The availability of large-scale datasets, such as Bengali Common Voice (Alam et al., 2022), OOD-Speech (Rakib et al., 2023a), and RegSpeech12 (Hassan et al., 2025), has created significant opportunities for advancing Bengali ASR systems. Continuous Bengali ASR has been benchmarked on the SHRUTI corpus using DNN-HMM and GMM-HMM-based models (Al Amin et al., 2019), achieving relatively low error rates. Leveraging CNN-RNN architectures, Saha et al. (2021) developed a gender- and speaker-independent ASR system. In Bangla-Wave (Rakib et al., 2023b), the integration of an n-gram language model is shown to yield notable performance improvements. Furthermore, comparative analysis demonstrates that the Wav2Vec-BERT model outperforms Whisper on the Bengali Common Voice dataset (Ridoy et al., 2025).
@@ C.1 Ben10 Dataset Analysis
The Ben10 dataset (Humayun et al., 2024) comprises speech recordings from 373 speakers across ten distinct regions of Bangladesh. The training set contains over 63 hours of audio sampled at 16 kHz. Table˜4 presents the region-wise distribution of audio files within the training set. Since the transcripts of the Ben10 test set are not publicly available, we instead employ the test set of the RegSpeech12 dataset (Hassan et al., 2025) for evaluation in our study.
@@ C.2 RegSpeech12 Dataset Analysis
The RegSpeech12 dataset (Hassan et al., 2025) is a spontaneous speech corpus encompassing twelve regional dialects of Bangladesh, comprising approximately 100 hours of speech data. In this study, we utilize the test split of the dataset, which contains around 10 hours of recordings. Table˜5 presents the region-wise distribution of the test set.
@@ Appendix F TTS Systems Evaluation
In this study, we conduct extensive experiments with three open-source TTS systems, all of which are based on the VITS architecture (Kim et al., 2021): MMS-TTS-Ben (Pratap et al., 2024), and the VITS-Bengali Male and Female variants (Hossen, 2023). To evaluate the quality of the synthesized speech, we employ the mean opinion score (MOS) metric.
For a robust comparison, we curate a diverse set of texts from the BanSpeech dataset (Samin et al., 2024), which contains audio-text pairs across thirteen categories. From each category, one representative text sample is selected, forming the test dataset, as detailed in §˜F.1. Each text sample is synthesized using all three TTS models. To assess naturalness and perceived audio quality, an experienced human rater with expertise in audio signal processing independently rates each generated audio on a 1 to 5 scale, where 1 indicates very poor quality and 5 represents excellent quality (Streijl et al., 2016).
@@ F.1 Text Samples from the BanSpeech Dataset
The BanSpeech dataset (Samin et al., 2024) comprises audio–text pairs spanning diverse categories. For evaluating the TTS systems, we use the text samples from this dataset, as illustrated in Fig.˜9.

### v2

@@ 4.1 Bengali Speech Assistants
Several speech assistant systems have been developed for the Bengali language in recent years. The ALAPI system (hasan2021alapi) introduces an open-domain Bengali conversational agent that processes recorded audio queries from users and generates response audio using AI techniques alongside a custom-built database. Shohojogi (arnab2023shohojogi), designed for the banking sector, provides voice-based customer support in Bengali. It integrates Wav2Vec2-based ASR (baevski2020wav2vec), query summarization, Google Text-to-Speech (gTTS), and a doc2vec model to retrieve relevant information for responses. Adrisya Sahayak (sultan2021adrisya) presents a desktop-based virtual speech assistant for visually impaired Bengali speakers, supporting computer operations, peripheral devices, and home appliance control.
In the healthcare domain, a voice-enabled Artificial Conversational Entity has been developed to automate service interactions in Bengali (pranto2021human). This system leverages a domain-specific database and similarity-matching strategies to generate responses to user queries. Extending beyond monolingual assistants, Disha (ullah2024disha) is a humanoid virtual assistant capable of interacting in both Bengali and English. It can address financial queries and perform real-time transactions. Beyond general-purpose applications, Bengali voice assistants have also been deployed in specialized domains such as providing information on metro rail services in Bangladesh (rahman2024comprehensive), assisting farmers with agricultural queries (divakar2021farmer), and offering accessible feminine healthcare support for marginalized women (puja2024exploring).
@@ 4.2 English Speech Assistants
Voice assistants such as Amazon Alexa, Apple Siri, Microsoft Cortana, and Google Assistant have become integral to modern human-computer interaction (lopez2017alexa). These systems, powered by artificial intelligence, are increasingly influencing daily life and societal practices (subhash2020artificial). Their development has been driven by advances in both signal processing and machine learning, which form the technological foundation of voice-based interaction (haeb2019speech).
Beyond their technical design, researchers have examined the broader impact of these systems. For example, flavian2023effects demonstrated that, compared to text-based recommendations, voice-based recommendations delivered by assistants have a stronger influence on consumer decision-making. User studies also indicate that voice assistants are most frequently employed for activities such as music, information search, and smart home (IoT) control (ammari2019music). Furthermore, their role has extended into education and healthcare contexts, where they support learning environments and assist older adults in managing everyday activities (terzopoulos2020voice; oewel2023voice).
@@ 4.3 Automatic Speech Recognition in Bengali
Research on Automatic Speech Recognition (ASR) in Bengali has been addressed through both system development and survey-driven studies (mridha2022study; tasnia2023overview; sultana2021recent). The availability of large-scale datasets, such as Bengali Common Voice (alam2022bengali), OOD-Speech (rakib2023ood), and RegSpeech12 (regspeech12), has created significant opportunities for advancing Bengali ASR systems. Continuous Bengali ASR has been benchmarked on the SHRUTI corpus using DNN-HMM and GMM-HMM-based models (al2019continuous), achieving relatively low error rates. Leveraging CNN-RNN architectures, saha2021development developed a gender- and speaker-independent ASR system. In Bangla-Wave (rakib2023bangla), the integration of an n-gram language model is shown to yield notable performance improvements. Furthermore, comparative analysis demonstrates that the Wav2Vec-BERT model outperforms Whisper on the Bengali Common Voice dataset (ridoy2025adaptability).
@@ C.1 Ben10 Dataset Analysis
The Ben10 dataset (ben10) comprises speech recordings from 373 speakers across ten distinct regions of Bangladesh. The training set contains over 63 hours of audio sampled at 16 kHz. Table˜4 presents the region-wise distribution of audio files within the training set. Since the transcripts of the Ben10 test set are not publicly available, we instead employ the test set of the RegSpeech12 dataset (regspeech12) for evaluation in our study.
@@ C.2 RegSpeech12 Dataset Analysis
The RegSpeech12 dataset (regspeech12) is a spontaneous speech corpus encompassing twelve regional dialects of Bangladesh, comprising approximately 100 hours of speech data. In this study, we utilize the test split of the dataset, which contains around 10 hours of recordings. Table˜5 presents the region-wise distribution of the test set.
@@ Appendix F TTS Systems Evaluation
In this study, we conduct extensive experiments with three open-source TTS systems, all of which are based on the VITS architecture (kim2021conditional): MMS-TTS-Ben (pratap2024scaling), and the VITS-Bengali Male and Female variants (ComprehensiveBanglaTTS2023). To evaluate the quality of the synthesized speech, we employ the mean opinion score (MOS) metric.
For a robust comparison, we curate a diverse set of texts from the BanSpeech dataset (samin2024banspeech), which contains audio-text pairs across thirteen categories. From each category, one representative text sample is selected, forming the test dataset, as detailed in §˜F.1. Each text sample is synthesized using all three TTS models. To assess naturalness and perceived audio quality, an experienced human rater with expertise in audio signal processing independently rates each generated audio on a 1 to 5 scale, where 1 indicates very poor quality and 5 represents excellent quality (streijl2016mean).
@@ F.1 Text Samples from the BanSpeech Dataset
The BanSpeech dataset (samin2024banspeech) comprises audio–text pairs spanning diverse categories. For evaluating the TTS systems, we use the text samples from this dataset, as illustrated in Fig.˜9.
@@ Appendix G End-to-End delay of BanglaTalk System
Appendix H User Study
To evaluate the user experience with the BanglaTalk system, a rigorous qualitative analysis is performed. Table˜9 presents a summary of the experimental results, while Fig.˜10 illustrates user interactions with the BanglaTalk system from two regions of Bangladesh – Sylhet and Mymensingh.
User Region Rating Mean $\uparrow$
Query_1 Query_2 Query_3 Query_4 Query_5
User_1 Sylhet 3.0 4.0 2.5 3.0 5 3.5
User_3 Sylhet 3.4 3.7 3.2 3.3 4.0 3.52
User_2 Mymensingh 2.0 4.0 3.8 4.2 3.3 3.46
User_4 Mymensingh 5.0 3.0 4.5 3.5 4.0 4.0
Mean 3.62
Table 9: User ratings on a 1–5 scale (where 1 indicates a poor experience and 5 indicates an excellent experience) for conversational interactions with the BanglaTalk speech assistant. Ratings are collected from four users representing two regional dialects – Sylhet and Mymensingh.
Figure 10: Conversational interactions with the BanglaTalk speech assistant using the regional dialects of Mymensingh (left) and Sylhet (right). Each example includes the user’s query, the transcript generated by the BRDialect ASR system, and the corresponding LLM-generated response. Although the WER of BRDialect is relatively high in some cases, the LLM effectively captures the context and produces appropriate responses.